OpenStudy (anonymous):

Solve the equation and check the solution(s). e^2x-2e^x-3=0

OpenStudy (raden):

let's assume that e^x = y, so it like quadratic : y^2 - 2y - 3 = 0 factor out then solve for y, first

OpenStudy (anonymous):

y=-1, y=3

OpenStudy (raden):

yes, that's right. now you have to resubtitute them to e^x = y. get e^x = -1 or e^x = 3 then solve for x

OpenStudy (ranga):

Since e^x can never be negative for any real value of x, you can discard e^x = -1. Solve e^x = 3. Take natural logarithm on both sides to solve for x.